OpenStudy (anonymous):

Louann and Carla received equal scores on a test made up of multiple choice questions and an essay. Louann got 14 multiple choice questions correct and received 24 points for her essay. Carla got 16 multiple choice questions correct and received 14 points for her essay. How many points was each multiple choice question worth?

OpenStudy (anonymous):

lets say each question multiple choice question was of x marks .. then louann marks = 14x + 24 carla marks = 16x + 14 now both are equal 14x + 24 = 16x + 14 10 = 2x x = 5 so marks for each MCQ (multiple choice question) was 5